OpenStudy (anonymous):

What is the slope of the line that passes through (-2, 7) and (4, 9)? A. slope = -1/3 B. slope = 1/3 C.slope = 3

OpenStudy (anonymous):

2/6 so simplify to 1/3. B.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ 7}})\quad &({\color{red}{ 4}}\quad ,&{\color{blue}{ 9}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

This what you need help with?

OpenStudy (anonymous):

9 - 7 and 4 - -2. That gives you 2/6. Divide both numerator and denominator by 2. 1/3.

OpenStudy (anonymous):

I did it the first time then i got A as the answer? @shopgirl1701

OpenStudy (anonymous):

Answer is B. You can also graph this really quick to make sure the line is rising (positive slope).

OpenStudy (anonymous):

Okay, here's where you went wrong. You probably said -4 - 2 = -6. But when you have a DOUBLE negative (4 MINUS negative 2), you get a positive. So 4 + 2 = 6.

OpenStudy (anonymous):

And you know 9 - 7 = 2. So 2/6 :)

OpenStudy (anonymous):

Thank you all!

OpenStudy (anonymous):

Sho thing.